OpenStudy (anonymous):

585*9%*6= 315.90 is this the correct answer for exact interest

OpenStudy (anonymous):

you mean simple interest? Never heard of exact interest i = prt

OpenStudy (anonymous):

no the exact interest

OpenStudy (anonymous):

P*r*t=I

OpenStudy (anonymous):

I just did the simple interest earlier

OpenStudy (anonymous):

585*0.09*6=315.9

OpenStudy (anonymous):

hank you very much

OpenStudy (anonymous):

exact interest is calculated based on the exact days on a 365-day calendar. so I am assuming that you meant 6 months 1/2 * 365= 182.5 days

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well actually, 6 mos = counting Jan - Jun then it is 181/365 that should be used for T